# Shape-Space Graphs: Fast and Collision-Free Path Planning for Soft Robots

### v1

@@ Abstract
Soft robots, inspired by elephant trunks or octopus arms, offer extraordinary flexibility to bend, twist, and elongate in ways that rigid robots cannot. However, their motion planning remains a challenge, especially in cluttered environments with obstacles, due to their highly nonlinear and infinite-dimensional kinematics. Here, we present a graph-based path planning tool for an elephant-trunk-inspired soft robotic arm designed with three artificial muscle fibers that allow for multimodal continuous deformation through contraction. Using a biomechanical model inspired by morphoelasticity and active filament theory, we precompute a shape library and construct a $k$-nearest neighbor graph in shape space, ensuring that each node corresponds to a mechanically accurate and physically valid robot shape. For the graph, we use signed distance functions to prune nodes and edges colliding with obstacles, and define multi-objective edge costs based on geometric distance and actuation effort, enabling energy-efficient planning with collision avoidance. We demonstrate that our algorithm reliably avoids obstacles and generates feasible paths within milliseconds from precomputed graphs using Dijkstra’s algorithm. We show that including energy costs can drastically reduce the actuation effort compared to geometry-only planning, at the expense of longer tip trajectories. Our results highlight the potential of shape-space graph search for fast and reliable path planning in the field of soft robotics, paving the way for real-time applications in surgical, industrial, and assistive settings.
@@ I Introduction
Soft continuum robots, inspired by biological structures such as elephant trunks, octopus arms, or plant stems, can bend, twist, and elongate in ways that rigid robots cannot [4]. This extreme flexibility enables them to navigate tight spaces, adapt to complex environments, and interact safely with delicate objects, making soft robotic structures promising for applications ranging from fruit harvesting [16, 47], search and rescue [7, 40] to minimally invasive surgery [46].
@@ I-A Related Work
Inverse kinematics-based methods. Many “conventional” path planning algorithms in rigid robotics rely on inverse kinematics: The desired robot or end-effector path is specified in $\mathcal{W}$-space. Then, at each point along the path, the inverse kinematics are solved to find the necessary configuration in $\mathcal{C}$-space, and these configurations are connected, possibly by interpolation, to form the continuous path. Such methods have been explored in soft robotics, but their highly complex kinematics make them prone to geometric constraint violations, local minima, non-uniqueness, infeasibility of the solution, or high computational cost in cluttered environments [38, 6, 13, 41, 3]. These drawbacks have motivated a shift towards methods that plan directly in $\mathcal{C}$-space, avoiding explicit inverse kinematics.
@@
Artificial potential methods. Here, the robot is treated as a moving point in $\mathcal{C}$-space that is drawn toward attractive potentials created by the goal configuration and repelled by obstacle potentials [22, 45]. The path emerges from following the gradient of this combined potential field. This approach has been validated computationally for curve-like obstacles [11] and experimentally for a cable-driven continuum robot [9], but since these methods are based on continuous optimization, they can suffer from local minima.
@@
Sampling-based methods. Algorithms such as rapidly exploring random trees (RRT) are well established in rigid robotics for their efficiency and simplicity when planning in $\mathcal{C}$-space [24]. Their extension RRT* adds asymptotic optimality [20]. For soft robots, the lack of closed-form inverse kinematics makes sampling challenging. Some works use relaxed kinematic constraints with collision checking [2]; others combine RRT* with additional constraints, such as safety optimization [31]. Special efforts have also been made in the field of concentric tube robots for minimally invasive surgery, where RRTs explore $\mathcal{C}$-space to find a path in $\mathcal{W}$-space [1, 48, 43, 26]. Alternatively, Jacobian-based inverse kinematics routines can project samples into feasible robot configurations [35].
@@
Graph-based methods. Also called roadmap techniques, these reduce the high-dimensional $\mathcal{C}$-space to a one-dimensional, searchable path on a graph. Graphs can be built by connecting nearby configurations of a precomputed library ($k$-nearest neighbors), by randomly sampling and connecting collision-free states as in probabilistic roadmaps [21], or by discretizing the space into a grid or motion primitives. Once built, the graph can be searched using algorithms such as Dijkstra [8] for optimality or A* [15] for faster heuristic search. Both allow custom edge costs such as geometric distance, actuation effort, or obstacle avoidance. In soft robotics, examples include a time-augmented $\mathcal{C}$-space graph for dynamic obstacle avoidance [33], grid discretization for a soft bending pneumatic actuator [42], nearest-neighbor graphs for concentric tube robots [39], and hybrid methods that generate configurations using inverse kinematics and then run Dijkstra’s search [34]. Most of these methods still operate in $\mathcal{C}$-space, even when the goal is specified in $\mathcal{W}$-space.
@@
Obstacle representation. Obstacles can be incorporated into path planning through a variety of methods, e.g., intrinsically in artificial potential fields [22], explicit geometric intersection checks [11], or volumetric occupancy maps. Among these, signed distance fields (SDFs) have become particularly popular for encoding the minimum distance from any point in the workspace to the nearest obstacle, enabling both fast collision detection and continuous clearance costs [23, 10, 28]. Although most SDFs are defined in $\mathcal{W}$-space, they are emerging in $\mathcal{C}$-space for high-dimensional planners [27, 29]. In soft robotics, SDF-based models of continuum robot shapes can accelerate collision checks even in cluttered environments [30], where repeated geometric queries are often a major computational bottleneck [11].
@@ I-B Contributions
To address the challenge of fast, collision-free path planning for soft robots, we propose a computationally efficient graph-based planning tool directly in $\mathcal{W}$-space, or more appropriately, shape space ($\mathcal{S}$-space), which is better suited for soft robots due to their continuum nature. Our method integrates a biomechanically exact forward kinematics model of a soft robotic arm [17] with a precomputed shape library and a $k$-nearest neighbor ($k$-NN) graph that guarantees that all nodes correspond to physically valid shapes, enabling multi-objective planning at high speed. By using signed distance fields (SDFs), clearance is precomputed and infeasible nodes are pruned before search, avoiding expensive collision checks during runtime. This combination of exact modeling and offline graph construction allows us to eliminate dependence on inverse kinematics during online planning and generate efficient and realistic paths. Our main contributions are as follows.
@@ II Biomechanical Soft Robot Model
  • slender: one dimension is much larger than the others,
@@
which is why we model it as an active filament. In contrast to passive filaments with fixed material and geometry properties, active filaments can actively contract, e.g., through magnetic actuation, biological growth, or muscular contraction. To describe such systems, we combine the theory of morphoelasticity [14, 37] with active filament theory [19].
@@
$\bm{r}^{\prime}(z)$ $=\zeta\bm{d}_{3},$ (1)
$\bm{d}_{i}^{\prime}(z)$ $=\zeta\bm{u}\times\bm{d}_{i},\quad i=\{1,2,3\},$ (2)
with the axial extension of the filament $\zeta$ and the Darboux curvature vector $\bm{u}=u_{1}\bm{d}_{1}+u_{2}\bm{d}_{2}+u_{3}\bm{d}_{3}$ describing the evolution of the director basis $\bm{D}$ along the filament [14, 37]. We denote the properties of the filament in the absence of external loading as intrinsic curvatures $\bm{u}=\bm{\hat{u}}$ and intrinsic extension $\zeta=\hat{\zeta}$.
@@
External loading. We define a deformed configuration $\mathcal{B}_{\text{d}}$ as a result of deforming the activated configuration $\mathcal{B}$ through external forces. We assume that the motion of the robotic arm is slow, and consider the quasi-static solution of the force and moment balance equations for morphoelastic rods [37, 18],
@@
$\bm{d}_{i}^{\prime}(z)$ $=\zeta\bm{u}\times\bm{d}_{i},\quad i=\{1,2,3\},$ (5b)
$\frac{\partial\bm{n}}{\partial z}+\hat{\zeta}\bm{f}$ $=\bm{0}$ (5c)
$\frac{\partial\bm{m}}{\partial z}+\frac{\partial\bm{r}}{\partial z}\times\bm{n}+\hat{\zeta}\bm{l}$ $=\bm{0}.$ (5d)
@@ III Shape-Space Graph Construction and Search
Unlike optimization-based strategies, library-based approaches provide a robust alternative for inverse problems [12, 32], as they avoid difficulties related to non-convex objective functions and local minima. Here, we adopt this concept in the context of path planning for the soft robot. The biomechanically inspired model is the baseline for generating a library of physically feasible robot shapes that are then used to build a graph in $\mathcal{S}$-space and find collision-free paths. A code example is provided on Github.
@@ III-C Collision Avoidance with Signed Distance Fields
such that $\phi(\mathbf{x})>0$ outside obstacles, $\phi(\mathbf{x})=0$ on their surface, and $\phi(\mathbf{x})<0$ inside. We here focus on relatively simple shapes, including cuboids and cylinders, for which solving the minimization problem $\min_{\bm{q}\in\partial\Omega}\|\bm{p}-\bm{q}\|$ is trivial and allows for a closed-form implementation. However, SDFs for more complex shapes may be implemented by solving the minimization problem numerically.
@@
that is, the total node clearance is the minimum of the node clearances of the individual obstacles. In practice, we precompute the node clearance $\{c^{(i)}\}$ offline using analytic obstacle SDFs, e.g., cylinders, boxes. We retain nodes with positive clearance $c>0$ and prune all nodes with $c\leq 0$. Finally, we perform edge sweeps by evaluating $\phi$ along linear interpolants between nodes, pruning edges that intersect obstacles.
@@ III-D Edge Weight Formulations
Assigning edge weights to the resulting candidate edges is a key feature of graph-based path planning, as this will define the shortest path, in terms of lowest cost, from a start node to an end node, and allows us to constrain certain paths and prefer others. With our shape library consisting of tuples $(\bm{\gamma}^{(i)},\bm{R}^{(i)})$ of activations and centerline shapes, we can formulate edge costs both in $\mathcal{A}$-space and in $\mathcal{S}$-space.
@@
Activation energy. To account for the effort of the liquid crystal elastomer fibers, we penalize large absolute activation levels through the edge weight, defined in $\mathcal{A}$-space,
@@
Activation rate. In addition to keeping the activation magnitude low, it is desirable to reduce rapid heating and cooling ramps in the liquid crystal elastomer fibers. Hence, we penalize large activation jumps between consecutive activation vectors,
@@
with weighting parameters $\alpha>0$, $\beta\geq 0$, $\delta\geq 0$. These parameters represent the trade-off between the different objectives.
@@ III-E Graph Search for Path Planning
where $i_{0}$ and $i_{m}$ denote the start and goal nodes and consecutive pairs $(i_{k},i_{k+1})$ are connected by edges $(i_{k},i_{k+1})\in E$.
@@ IV Results
We evaluate the proposed planning framework with the precomputed shape library of the soft robotic arm that contains $N=100,000$ samples with a spatial discretization of $n_{z}=100$. We build the $k$-NN graph with $k=20$ neighbors. Here we evaluate our path-finding tool purely computationally, but note that it can be validated experimentally with our three-fiber prototype [25].
@@
Summary metrics for both scenarios are presented in Table I.
@@ V Discussion
This work introduces a graph-based path planning tool for soft robotic arms that ensures physically valid configurations by coupling a biomechanically accurate model with a precomputed shape-space graph. The $k$-NN graph, constructed with $k=20$ neighbors from a shape library of 100,000 samples and multi-objective edge costs, quickly identifies collision-free, energy-efficient paths. Our results suggest that incorporating energy costs naturally leads to trajectories that trade longer paths for reduced activation effort, because elongated fibers are less energy-demanding. The weighing parameters encoded in the edge cost can be adjusted flexibly and tailored to specific applications and priorities. Scalability remains a key challenge: We found that 100,000 samples provide a good trade-off, while larger shape libraries can improve trajectory smoothness and reduce oscillations in activation sequences. However, larger libraries increase memory and computational costs. Potential remedies include compressing the shape library via encoder-decoder architectures to balance smoothness and efficiency.

### v2

@@ Abstract
Soft robots, inspired by elephant trunks or octopus arms, offer extraordinary flexibility to bend, twist, and elongate in ways that rigid robots cannot. However, their motion planning remains a challenge, especially in cluttered environments with obstacles, due to their highly nonlinear and infinite-dimensional kinematics. Here, we present a graph-based path planning tool for an elephant-trunk-inspired soft robot designed with three artificial muscle fibers that allow for continuous deformation through contraction. Using a biomechanical model that integrates morphoelastic and active filament theories, we precompute a shape library and construct a $k$-nearest neighbor graph in shape space, ensuring that each node corresponds to a valid robot shape. For the graph, we use signed distance functions to prune nodes and edges colliding with obstacles, and define multi-objective edge costs based on geometric distance and actuation effort, enabling energy-aware planning with collision avoidance. We demonstrate that our algorithm reliably avoids obstacles and generates feasible paths within milliseconds from precomputed graphs using Dijkstra’s algorithm. We show that including energy costs can drastically reduce the actuation effort compared to geometry-only planning, at the expense of longer tip trajectories. Our results highlight the potential of shape-space graph search for fast and reliable path planning in the field of soft robotics, paving the way for real-time applications in surgical, industrial, and assistive settings.
@@ I Introduction
Soft continuum robots, inspired by biological structures such as elephant trunks, octopus arms, or plant stems, can bend, twist, and elongate in ways that rigid robots cannot. This extreme flexibility enables them to navigate tight spaces, adapt to complex environments, and interact safely with delicate objects, making soft robotic structures promising for applications ranging from fruit harvesting [33], search and rescue [4] to minimally invasive surgery [32].
@@ I-A Related Work
Inverse kinematics-based methods. Many “conventional” path planning algorithms in rigid robotics rely on inverse kinematics: The desired robot or end-effector path is specified in $\mathcal{W}$-space. Then, at each point along the path, the inverse kinematics are solved to find the necessary configuration in $\mathcal{C}$-space, and these configurations are connected, possibly by interpolation, to form the continuous path. Such methods have been explored in soft robotics, but their highly complex kinematics make them prone to geometric constraint violations, local minima, non-uniqueness, infeasibility of the solution, or high computational cost in cluttered environments [28, 3, 9]. These drawbacks have motivated a shift towards methods that plan directly in $\mathcal{C}$-space, avoiding explicit inverse kinematics.
@@
Artificial potential methods. Here, the robot is treated as a moving point in $\mathcal{C}$-space that is drawn toward attractive potentials created by the goal configuration and repelled by obstacle potentials [14]. The path emerges from following the gradient of this combined potential field. This approach has been validated computationally for curve-like obstacles [7] and experimentally for a cable-driven continuum robot [6], but since these methods are based on continuous optimization, they can suffer from local minima.
@@
Sampling-based methods. Algorithms such as rapidly exploring random trees (RRT) are well established in rigid robotics for their efficiency and simplicity when planning in $\mathcal{C}$-space [16]. Their extension RRT* adds asymptotic optimality [12]. For soft robots, the lack of closed-form inverse kinematics makes sampling challenging. Some works use relaxed kinematic constraints with collision checking [2]; others combine RRT* with additional constraints, such as safety optimization [23], or consider multitarget paths [35]. Special efforts have also been made in the field of concentric tube robots for minimally invasive surgery, where RRTs explore $\mathcal{C}$-space to find a path in $\mathcal{W}$-space [1, 34, 18]. Alternatively, Jacobian-based inverse kinematics routines can project samples into feasible robot configurations [26].
@@
Graph-based methods. Also called roadmap techniques, these reduce the high-dimensional $\mathcal{C}$-space to a one-dimensional, searchable path on a graph. Graphs can be built by connecting nearby configurations of a precomputed library ($k$-nearest neighbors), by randomly sampling and connecting collision-free states as in probabilistic roadmaps [13], or by discretizing the space into a grid or motion primitives. Once built, the graph can be searched using algorithms such as Dijkstra [5]. It allows for custom edge costs such as geometric distance, actuation effort, or obstacle avoidance. In soft robotics, examples include a time-augmented $\mathcal{C}$-space graph for dynamic obstacle avoidance [24], grid discretization for a soft bending pneumatic actuator [31], nearest-neighbor graphs for concentric tube robots [29], and hybrid methods that generate configurations using inverse kinematics and then run Dijkstra’s search [25]. Most of these methods still operate in $\mathcal{C}$-space, even when the goal is specified in $\mathcal{W}$-space.
@@
Obstacle representation. Obstacles can be incorporated into path planning through a variety of methods, e.g., intrinsically in artificial potential fields [14], explicit geometric intersection checks [7], or volumetric occupancy maps. Among these, signed distance fields (SDFs) have become particularly popular for encoding the minimum distance from any point in the workspace to the nearest obstacle, enabling both fast collision detection and continuous clearance costs [15, 20]. Although most SDFs are defined in $\mathcal{W}$-space, they are emerging in $\mathcal{C}$-space for high-dimensional planners [19, 21]. In soft robotics, SDF-based models of continuum robot shapes can accelerate collision checks even in cluttered environments [22], where repeated geometric queries are often a major computational bottleneck [7].
@@ I-B Contributions
To address the challenge of fast, collision-free path planning for soft robots, we propose a computationally efficient graph-based planning tool directly in $\mathcal{W}$-space, or more appropriately, shape space ($\mathcal{S}$-space), which is better suited for soft robots due to their continuum nature. Our method integrates a biomechanically exact forward kinematics model of a soft robotic arm [10] with a precomputed shape library and a $k$-nearest neighbor ($k$-NN) graph that guarantees that all nodes correspond to physically valid shapes, enabling multi-objective planning at high speed. By using signed distance fields (SDFs), clearance is precomputed and infeasible nodes are pruned before search, avoiding expensive collision checks during runtime. This combination of exact modeling and offline graph construction allows us to eliminate dependence on inverse kinematics during planning and generate efficient and realistic paths. Our main contributions are as follows.
@@ II Biomechanical Soft Robot Model
  • slender: it is orders of magnitude longer than it is thick,
@@
which is why we model it as an active filament. In contrast to passive filaments with fixed material and geometry properties, active filaments can actively contract, e.g., through magnetic actuation, biological growth, or muscular contraction. To describe such systems, we use the theory of morphoelasticity [27, 11]. We note that Cosserat-rod based formulations play a central role in soft robotics that offer geometrically exact rod models and reduced-order strain parameterizations [30]. The theory of morphoelasticity extends classical Cosserat-rod formulations by allowing for contraction and growth.
@@
$\bm{r}^{\prime}(z)$ $=\zeta\bm{d}_{3},$ (1a)
$\bm{d}_{i}^{\prime}(z)$ $=\zeta\bm{u}\times\bm{d}_{i},\quad i=\{1,2,3\},$ (1b)
with the axial extension of the filament $\zeta$ and the Darboux curvature vector $\bm{u}=u_{1}\bm{d}_{1}+u_{2}\bm{d}_{2}+u_{3}\bm{d}_{3}$ describing the evolution of the director basis $\bm{D}$ along the filament [27]. We denote the properties of the filament in the absence of external loading as intrinsic curvatures $\bm{u}=\bm{\hat{u}}$ and intrinsic extension $\zeta=\hat{\zeta}$, i.e., the intrinsic shape is the shape induced by activation in the absence of external loads.
@@
External loading. We define a deformed configuration $\mathcal{B}_{\text{d}}$ as a result of deforming the activated configuration $\mathcal{B}$ through external forces. We assume that the motion of the robotic arm is slow, and consider the quasi-static solution of the force and moment balance equations for morphoelastic rods [27],
@@
$\bm{d}_{i}^{\prime}(z)$ $=\hat{\zeta}\bm{u}\times\bm{d}_{i},\quad i=\{1,2,3\},$ (4b)
$\bm{0}$ $=\frac{\partial\bm{n}}{\partial z}+\hat{\zeta}\bm{f}$ (4c)
$\bm{0}$ $=\frac{\partial\bm{m}}{\partial z}+\frac{\partial\bm{r}}{\partial z}\times\bm{n}+\hat{\zeta}\bm{l}.$ (4d)
@@ III Shape-Space Graph Construction and Search
Unlike optimization-based strategies, library-based approaches provide a robust alternative for inverse problems [8], as they avoid difficulties related to non-convex objective functions and local minima. Here, we adopt this concept in the context of path planning for the soft robot. The biomechanically inspired model is the baseline for generating a library of physically feasible robot shapes that are then used to build a graph in $\mathcal{S}$-space and find collision-free paths. A code example is provided on Github.
@@ III-C Collision Avoidance with Signed Distance Fields
such that $\phi(\mathbf{p})>0$ outside obstacles, $\phi(\mathbf{p})=0$ on their surface, and $\phi(\mathbf{p})<0$ inside. We here focus on relatively simple shapes, including cuboids and cylinders, for which solving the minimization problem $\min_{\bm{q}\in\partial\Omega}\|\bm{p}-\bm{q}\|$ is trivial and allows for a closed-form implementation. However, SDFs for more complex shapes may be implemented by solving the minimization problem numerically.
@@
that is, the total node clearance is the minimum of the node clearances of the individual obstacles. In practice, we precompute the node clearance $\{c^{(i)}\}$ offline using analytic obstacle SDFs, e.g., cylinders, boxes. We retain nodes with positive clearance $c>0$ and prune all nodes with $c\leq 0$. Finally, we perform edge sweeps by evaluating $\phi$ along linear interpolants between nodes, pruning edges that intersect obstacles. While we focus on simple analytical obstacles, the SDF-based pruning is not restricted to simple geometries, but works for any obstacle for which we can evaluate an SDF, including non-convex or mesh-based obstacles.
@@ III-D Edge Weight Formulations
Assigning edge weights to the resulting candidate edges is a key feature of graph-based path planning, as this will define the shortest path, in terms of lowest cost, from a start node to an end node. It also allows us to constrain certain paths and prefer others. With our shape library consisting of tuples $(\bm{\gamma}^{(i)},\bm{R}^{(i)})$ of activations and centerline shapes, we can formulate edge costs both in $\mathcal{A}$-space and in $\mathcal{S}$-space.
@@
Activation energy. To account for the effort of the actuators, we penalize large absolute activation levels through the edge weight, defined in $\mathcal{A}$-space,
@@
Activation rate. In addition to keeping the activation magnitude low, it is desirable to reduce rapid heating and cooling ramps for actuators such as liquid crystal elastomer fibers. Hence, we penalize large activation jumps between consecutive activation vectors,
@@
with weighting parameters $\alpha>0$, $\beta\geq 0$, $\delta\geq 0$. They represent the trade-off between the different objectives.
@@ III-E Graph Search for Path Planning
where $i_{0}$ and $i_{m}$ denote the start and goal nodes and consecutive pairs $(i_{k},i_{k+1})$ are connected by edges $(i_{k},i_{k+1})\in E$. To obtain a continuous actuation over time, we can interpolate within these discrete sequences.
@@ IV Results
[…]
@@
[…]
@@ IV-B Multi-Waypoint Route Through Cylindrical Obstacles
In the second scenario, the soft robot navigates between two cylindrical obstacles and reaches multiple shapes along the route, beginning and ending in the zero-position. Such scenarios are motivated by grippers needing to place objects in various locations while avoiding obstacles like shelves or machinery. We select waypoints as shapes from the precomputed library, place them close to obstacles to increase difficulty, and plan the motion as a sequence of local start–goal problems between two consecutive waypoints.
@@ V Discussion
This work introduces a graph-based path planning tool for soft robotic arms that ensures physically valid configurations by coupling a biomechanical model with a precomputed shape-space graph. The $k$-NN graph, constructed with $k=20$ neighbors from a shape library of 100,000 samples and multi-objective edge costs, quickly identifies collision-free, energy-aware paths. Our results suggest that incorporating energy costs naturally leads to trajectories that trade longer paths for reduced activation effort, because elongated fibers are less energy-demanding.
Limitations. Scalability remains a challenge: 100,000100{,}000 samples provide a good trade-off between smoothness and memory/computation time, and larger libraries may need compression techniques such as encoder–decoder architectures to remain practical. Still, building the graph takes up to several minutes, limiting the approach to static or infrequently changing obstacles. For more dynamic environments, incremental updates of the graph need to be explored. To include task-specific applications, the edge costs can be augmented with penalties on deviations from desired end-effector paths in workspace. Generally, a systematic exploration of parameters associated with generating the graph and costs is necessary. Finally, all results are computational, and validation on hardware is an important next step. Limitations coupled to the experimental validation include that (i) the method plans quasi-static sequences without enforcing dynamic feasibility or actuator-rate constraints, so continuous execution on hardware might require coupling to an appropriate dynamic controller, and that (ii) the shape library is generated for a fixed, history-independent activation–shape relation under given loading: cyclic effects and hysteresis in real actuators may shift the reachable shapes over time and would then require updating or adapting the library. We envision using the same three-fiber LCE prototype as developed previously [17], where each fiber contains an embedded resistive wire and has its own electrical channel, so the three fibers can be Joule-heated and controlled independently to generate different deformation modes. Alternatively, a similar fiber architecture could be realized with tendon- or cable-driven robots, where cables routed in the same longitudinal and helical layout are actuated by motors to reproduce the same bending and twisting capabilities.